# Long-range Meta-path Search on Large-scale Heterogeneous Graphs

### v4

@@ 6.1 Datasets and Baselines
We evaluate LMSPS on the representative large-scale dataset ogbn-mag from OGB challenge (Hu et al., 2021), and three widely-used heterogeneous graphs including DBLP, IMDB, and ACM from HGB benchmark (Lv et al., 2021) on the node classification task. Please see Appendix D.1 for details of all datasets. For the ogbn-mag dataset, the results are compared to the OGB leaderboard, and all scores are the average of $10$ separate training. For the three datasets from HGB, the results are compared to the baseline scores from the HGB paper and several recent advanced methods. Following the convention (Lv et al., 2021; Yang et al., 2023), all scores are the average of $5$ separate local data partitions. GEMS (Han et al., 2020) and RMS-HRec (Ning et al., 2022) are ignored because they are designed for recommendation tasks. RL-HGNN (Zhong et al., 2020) is omitted due to a lack of source code. HINormer (Mao et al., 2023) also doesn’t provide source code but uses two datasets from HGB, we show its reported results.
6.2 Training Details
We set the number of selected meta-paths $M=30$ for all datasets. The final search space $V=60$. The maximum hop is $6$ for ogbn-mag, DBLP and $5$ for IMDB, ACM. All $K$ architecture parameters $\alpha_{1},\alpha_{2},\cdots,\alpha_{K}$ are initialized as $1$s. For searching in the super-net, we train for $200$ epochs. To train the target-net, we use an early stop mechanism based on the validation performance to promise full training. See Appendix D.2 for more details.
6.3 Performance Comparison
Results on Ogbn-mag. To answer Q1, we report the test accuracy of our proposed LMSPS and baselines in Table 3. Random means the result of replacing our searched meta-paths with $30$ random meta-paths. We show the average result of 20 random samples. Because ogbn-mag is much more challenging than other datasets, existing methods usually use extra tricks. label indicates a kind of data augmentation utilizing labels as extra inputs to provide enhancements (Wang & Leskovec, 2020; Shi et al., 2021; Yang et al., 2023). ms means using multi-stage learning  (Li et al., 2018; Sun et al., 2020). It can be seen that LMSPS achieves state-of-the-art performance under each condition. For example, LMSPS outperforms the SOTA method SeHGNN by a large margin of $3.04\%$. The advantage of LMSPS compared to SeHGNN mainly comes from the unique ability to selectively leverage effective heterogeneous information under larger receptive fields, indicating the necessity of long-range meta-paths. SeHGNN shows the results with extra embeddings as an enhancement in the original paper. Though we cannot compare results with this trick due to the untouchability of its embedding file, LMSPS still outperforms their best result.
Performance on HGB Benchmark. Table 3 shows the Micro-F1 scores of LMSPS and baselines. The 1st, 2nd, and 3rd blocks are metapath-free, metapath-based, and NAS-based methods, respectively. The 4th block is our method. Every block contains a method yielding highly competitive performance, indicating that the HGNN field is very active and each mechanism has its advantage. Nevertheless, LMSPS can still consistently perform best on all three datasets.
6.4 Necessity of Long-range Dependency
To answer Q2 and explore the necessity of long-range dependency on heterogeneous graphs, we construct four large-scale datasets with high sparsity based on ogbn-mag. To avoid inappropriate preference seed settings of randomly removing, we construct fixed heterogeneous graphs by limiting the maximum in-degree related to edge type. Specifically, we gradually reduce the maximum in-degree related to edge type in ogbn-mag from $50$ to $5$ but leave all nodes unchanged. Details of the four datasets are listed in Appendix D.1. The test accuracy of LMSPS and SOTA method SeHGNN are shown in Table 4. LMSPS outperforms SeHGNN more significantly with the increasing sparsity. In addition, the leading gap of LMSPS over SeHGNN is more than $4.7\%$ on the highly sparse dataset ogbn-mag-5. The main difference between SeHGNN and LMSPS is that the former cannot utilize large hops and only use hop $2$ while the latter has a maximum hop of $6$, demonstrating that long-range dependencies are more effective with increased sparsity and decreased direct neighbors on heterogeneous graphs.
@@ 6.4 Necessity of Long-range Dependency
(a) Micro-F1
(b) GPU memory cost
(c) Time cost
6.5 Analysis on Large Maximum Hops
To answer Q3, we conduct experiments to compare the performance, memory, and efficiency of LMSPS with the best metapath-free method with source code, Simple-HGN (Lv et al., 2021), and the best metapath-based method, SeHGNN (Yang et al., 2023), on DBLP. Figure 3 (a) shows that LMSPS has consistent performance with the increment of maximum hop. The failure of Simple-HGN demonstrates its attention mechanism can not overcome the over-smoothing issue or eliminate the effects of noise under large hop. Figure 3 (b), (c) illustrate each training epoch’s average memory and time costs relative to the maximum hop or layer. We can observe that the consumption of SeHGNN exponentially grows, and the consumption of Simple-HGN linearly increases, which is consistent with their time complexity as listed in Table 1. At the same time, LMSPS has almost constant consumption as the maximum hop grows. Figure 3 (c) shows the time cost of LMSPS in the search stage, which also approximates a constant when the number of meta-paths is larger than $M=30$.
Table 5: Experiments on ogbn-mag to analyze the performance of SeHGNN and LMSPS under different maximum hops. #MP is the number of corresponding meta-paths of different maximum hops. “$OOM$” means out of memory.
Max hop SeHGNN LMSPS
(#MP) Time (s) Test accuracy Time (s) Test accuracy
$1$ $(4)$ 4.354.35~{}~{}4.354.35 $47.18\pm 0.28$ 3.983.98~{}~{}3.983.98 $46.88\pm 0.10$
$2$ $(10)$ 6.446.44~{}~{}6.446.44 $51.79\pm 0.24$ 5.635.63~{}~{}5.635.63 $51.91\pm 0.13$
$3$ $(23)$ $11.28$ $52.44\pm 0.16$ $10.02$ $52.72\pm 0.24$
$4$ $(50)$ $OOM$ $OOM$ $14.34$ $53.43\pm 0.18$
$5$ $(107)$ $OOM$ $OOM$ $14.77$ $53.90\pm 0.19$
$6$ $(226)$ $OOM$ $OOM$ $14.71$ $\textbf{54.83}\pm\textbf{0.20}$
We also compare the performance and training time of LMSPS with SeHGNN under different maximum hops on large-scale dataset ogbn-mag. When the maximum hop $l=1,2,3$, we utilize the full meta-paths set because the number of target-node-related meta-paths is smaller than $M=30$. Following the convention (Lv et al., 2021; Yang et al., 2023), we measure the average time consumption of one epoch for each model. As shown in Table 5, the performance of LMSPS keeps increases as the maximum hop value grows. It indicates LMSPS can overcome the issues caused by utilizing long-range dependency, e.g., over-smoothing and noise. In addition, when the number of meta-paths is larger than $30$, the training time of LMSPS is stable under different maximum hops. More results about efficiency are shown in Appendix D.4.
6.6 Generalization of the Searched Meta-paths
The effective meta-paths mainly depend on the dataset and task instead of the architecture. To demonstrate the effectiveness of searched meta-paths, on the one hand, the meta-paths should be effective in the proposed model. On the other hand, the meta-paths should be effective after being generalized to other HGNNs. Because finding meta-paths that work effectively across various HGNNs is a tough task, it has not been achieved by previous works. To answer Q4, we verify the generalization of our searched meta-path on the most famous HGNN HAN (Wang et al., 2019) and the SOTA method SeHGNN (Yang et al., 2023). The Micro-F1 scores on three representative datasets are shown in Table 6. After simply replacing the original meta-path set with our searched meta-paths and keeping other settings unchanged, the performance of both methods improves rather than decreases, demonstrating the effectiveness of our searched meta-paths. In addition, we have shown the interpretability searched meta-paths in Appendix C and explored the effectiveness of our search algorithm in Appendix D.3.
Method DBLP IMDB ACM
HAN $92.05\pm 0.62$ $64.63\pm 0.58$ $90.79\pm 0.43$
HAN* $93.54\pm 0.15$ $65.89\pm 0.52$ $92.28\pm 0.47$
SeHGNN $95.24\pm 0.13$ $68.21\pm 0.32$ $93.87\pm 0.50$
SeHGNN* $95.57\pm 0.23$ $68.59\pm 0.24$ $94.46\pm 0.18$

### v6

@@ 6.1 Datasets and Baselines
We evaluate LMSPS on several representative heterogeneous graph datasets, including DBLP, IMDB, ACM, and Freebase from HGB benchmark [32], and the large-scale dataset OGBN-MAG from OGB challenge [18]. The statistics of the datasets are listed in Table 7. The details about all datasets, baselines, and experiment settings are recorded in Appendix A and B.
Table 1: Performance on small and large datasets. Best is in bold, and the runner-up is underlined.
[b] Method DBLP IMDB ACM Freebase OGBN-MAG Macro-F1 Micro-F1 Macro-F1 Micro-F1 Macro-F1 Micro-F1 Macro-F1 Micro-F1 Test Acc. MLP [17] - - - - - - - - $26.92\pm 0.26$ GraphSAGE [13] - - - - - - - - $46.78\pm 0.67$ RGCN [39] $91.52\pm 0.50$ $92.07\pm 0.50$ $58.85\pm 0.26$ $62.05\pm 0.15$ $91.55\pm 0.74$ $91.41\pm 0.75$ $46.78\pm 0.77$ $58.33\pm 1.57$ $47.37\pm 0.48$ HAN [46] $91.67\pm 0.49$ $92.05\pm 0.62$ $57.74\pm 0.96$ $64.63\pm 0.58$ $90.89\pm 0.43$ $90.79\pm 0.43$ $21.31\pm 1.68$ $54.77\pm 1.40$ $OOM$ GTN [55] $93.52\pm 0.55$ $93.97\pm 0.54$ $60.47\pm 0.98$ $65.14\pm 0.45$ $91.31\pm 0.70$ $91.20\pm 0.71$ $OOM$ $OOM$ $OOM$ RSHN [62] $93.34\pm 0.58$ $93.81\pm 0.55$ $59.85\pm 3.21$ $64.22\pm 1.03$ $90.50\pm 1.51$ $90.32\pm 1.54$ $OOM$ $OOM$ $OOM$ HetGNN [56] $91.76\pm 0.43$ $92.33\pm 0.41$ $48.25\pm 0.67$ $51.16\pm 0.65$ $85.91\pm 0.25$ $86.05\pm 0.25$ $OOM$ $OOM$ $OOM$ MAGNN [10] $93.28\pm 0.51$ $93.76\pm 0.45$ $56.49\pm 3.20$ $64.67\pm 1.67$ $90.88\pm 0.64$ $90.77\pm 0.65$ $OOM$ $OOM$ $OOM$ HetSANN [16] $78.55\pm 2.42$ $80.56\pm 1.50$ $49.47\pm 1.21$ $57.68\pm 0.44$ $90.02\pm 0.35$ $89.91\pm 0.37$ $OOM$ $OOM$ $OOM$ GCN [24] $90.84\pm 0.32$ $91.47\pm 0.34$ $57.88\pm 1.18$ $64.82\pm 0.64$ $92.17\pm 0.24$ $92.12\pm 0.23$ $27.84\pm 3.13$ $60.23\pm 0.92$ $OOM$ GAT [44] $93.83\pm 0.27$ $93.39\pm 0.30$ $58.94\pm 1.35$ $64.86\pm 0.43$ $92.26\pm 0.94$ $92.19\pm 0.93$ $40.74\pm 2.58$ $65.26\pm 0.80$ $OOM$ Simple-HGN [32] $94.01\pm 0.24$ $94.46\pm 0.22$ $63.53\pm 1.36$ $67.36\pm 0.57$ $93.42\pm 0.44$ $93.35\pm 0.45$ $47.72\pm 1.48$ $66.29\pm 0.45$ $OOM$ HGT [19] $93.01\pm 0.23$ $93.49\pm 0.25$ $63.00\pm 1.19$ $67.20\pm 0.57$ $91.12\pm 0.76$ $91.00\pm 0.76$ $29.28\pm 2.52$ $60.51\pm 1.16$ $46.78\pm 0.42$ GraphMSE [29] $94.08\pm 0.14$ $94.44\pm 0.13$ $57.60\pm 2.13$ $62.37\pm 1.03$ $92.58\pm 0.50$ $92.54\pm 0.14$ $OOM$ $OOM$ $OOM$ DiffMG [7] $94.01\pm 0.37$ $94.20\pm 0.36$ $58.09\pm 1.35$ $59.75\pm 1.23$ $88.16\pm 2.83$ $88.07\pm 3.04$ $OOM$ $OOM$ $OOM$ Random $93.57\pm 0.64$ $93.84\pm 0.53$ $52.13\pm 0.74$ $53.83\pm 0.66$ $90.91\pm 1.02$ $90.82\pm 0.93$ $21.22\pm 2.58$ $37.54\pm 2.66$ $35.14\pm 3.78$ NARS [54] $94.18\pm 0.47$ $94.61\pm 0.42$ $63.51\pm 0.68$ $66.18\pm 0.70$ $93.36\pm 0.32$ $93.31\pm 0.33$ $49.98\pm 1.77$ $63.26\pm 1.26$ $50.66\pm 0.22$ space4HGNN [58] $94.24\pm 0.42$ $94.63\pm 0.40$ $61.57\pm 1.19$ $63.96\pm 0.43$ $92.50\pm 0.14$ $92.38\pm 0.10$ $41.37\pm 4.49$ $65.66\pm 4.94$ $OOM$ PMMM [25] $94.82\pm 0.26$ $95.14\pm 0.22$ $65.81\pm 0.29$ $67.58\pm 0.22$ $93.78\pm 0.25$ $93.71\pm 0.17$ $OOM$ $OOM$ $OOM$ HINormer [34] $94.57\pm 0.23$ $94.94\pm 0.21$ $64.65\pm 0.53$ $67.83\pm 0.34$ $93.91\pm 0.42$ $93.83\pm 0.45$ $\underline{52.18\pm 0.39}$ $64.92\pm 0.43$ $OOM$ SeHGNN [51] $94.86\pm 0.14$ $95.24\pm 0.13$ $\underline{66.63\pm 0.34}$ $68.21\pm 0.32$ $93.95\pm 0.48$ $93.87\pm 0.50$ $50.71\pm 0.44$ $63.41\pm 0.47$ $51.45\pm 0.29$ SlotGAT [61] $94.95\pm 0.20$ $\underline{95.31\pm 0.19}$ $64.05\pm 0.60$ $\underline{68.64\pm 0.33}$ $\underline{93.99\pm 0.23}$ $\underline{94.06\pm 0.22}$ $49.68\pm 1.97$ $\textbf{66.83}\pm\textbf{0.30}$ $OOM$ LMSPS (ours) $\textbf{95.35}\pm\textbf{0.22}$ $\textbf{95.66}\pm\textbf{0.20}$ $\textbf{66.99}\pm\textbf{0.32}$ $\textbf{68.70}\pm\textbf{0.26}$ $\textbf{94.73}\pm\textbf{0.41}$ $\textbf{94.69}\pm\textbf{0.36}$ $\textbf{53.26}\pm\textbf{0.47}$ $\underline{66.09\pm 0.51}$ $\textbf{54.83}\pm\textbf{0.20}$
    OGBN-MAG is a large dataset with nodes’ numbers 10 to 175 times that of the other four datasets.
6.2 Performance Analysis
To answer Q1, we report the performance of our approaches and baselines in Table 1. Random means the result of replacing our searched meta-paths with $30$ random meta-paths. We show the average result of 20 random samples. Based on the results, we have the following observations. First, LMSPS outperforms all baselines for different metrics on almost all datasets except Micro-F1 scores on Freebase, sometimes by a significant margin, which validates the power of LMSPS. For example, on the largest dataset OGBN-MAG, LMSPS achieves $54.83\%$ test accuracy, while the best competitor has $51.45\%$ test accuracy. Second, all metapath-free methods encounter out-of-memory (OOM) issues when dealing with large datasets, including highly competitive methods, HINormer and SlotGAT, indicating the advantage of employing meta-paths electively aggregating neighbors on the meta-path instances. Third, although LMSPS is an MLP-based method, the pure MLP method [17] has the worst performance with only $26.92\%$ test accuracy on OGBN-MAG, validating the advantages of pre-processing and meta-path search. Finally, LMSPS outperforms Random significantly, further indicating the importance of meta-path search. More comparisons with top method combinations on OGBN-MAG are shown in Table 10 of Appendix E.2, in which LMSPS also shows the best performance.
6.3 Analysis on Large Maximum Hops
To answer Q2, we conduct experiments to compare the performance, memory, and efficiency of LMSPS with the method of HGB benchmark, Simple-HGN [32], and the best metapath-based method, SeHGNN [51], on DBLP. Figure 3 (a) shows that LMSPS has consistent performance with the increment of maximum hop. The failure of Simple-HGN demonstrates its attention mechanism can not overcome the over-smoothing issue or eliminate the effects of noise under large hop. Figure 3 (b), (c) illustrate each training epoch’s average memory and time costs relative to the maximum hop or layer. We can observe that the consumption of SeHGNN exponentially grows, and the consumption of Simple-HGN linearly increases, which is consistent with their time complexity as listed in Table 8. Meanwhile, LMSPS has almost constant consumption as the maximum hop grows. Figure 3 (c) shows the time cost of LMSPS in the search stage, which also approximates a constant when the number of meta-paths is larger than $M=30$. More results about efficiency are shown in Figure 4 of Appendix E.1.
(a) Micro-F1
(b) GPU memory cost
(c) Time cost
We also compare the performance and training time of LMSPS with SeHGNN under different maximum hops on large-scale dataset OGBN-MAG. When the maximum hop $l=1,2,3$, we utilize the full meta-path set because the number of meta-paths is smaller than $M=30$. Following the convention [32, 51], we measure the average time consumption of one epoch for each model. As shown in Table 3, the performance of LMSPS keeps increasing as the maximum hop value grows. It indicates that LMSPS can overcome the issues caused by utilizing long-range dependency, e.g., over-smoothing and noise. In addition, when the number of meta-paths is larger than $30$, the training time of LMSPS is stable under different maximum hops.
6.4 Effectiveness of the Search Algorithm and Searched Meta-paths
Table 2: Experiments on OGBN-MAG to analyze the performance of SeHGNN and LMSPS under different maximum hops. #MP is the number of meta-paths under different maximum hops.
Max hop #MP SeHGNN LMSPS
Time Test accuracy Time Test accuracy
1 4   4.35 $47.18\pm 0.28$   3.98 $46.88\pm 0.10$
2 10   6.44 $51.79\pm 0.24$   5.63 $51.91\pm 0.13$
3 23 11.28 $52.44\pm 0.16$ 10.02 $52.72\pm 0.24$
4 50 OOM OOM 14.34 $53.43\pm 0.18$
5 107 OOM OOM 14.77 $53.90\pm 0.19$
6 226 OOM OOM 14.71 $\textbf{54.83}\pm\textbf{0.20}$
Method DBLP IMDB ACM Freebase
HAN $95.44\pm 0.14$ $65.95\pm 0.31$ $90.66\pm 0.30$
GTN $95.33\pm 0.05$ $65.99\pm 0.16$ $90.66\pm 0.30$
DARTS $95.35\pm 0.17$ $66.23\pm 0.14$ $93.45\pm 0.13$ $63.25\pm 0.42$
SPOS $95.41\pm 0.43$ $67.10\pm 0.29$ $93.64\pm 0.37$ $64.02\pm 0.62$
DiffMG $95.45\pm 0.49$ $66.98\pm 0.37$ $93.61\pm 0.45$ $64.56\pm 0.78$
PMMM $95.48\pm 0.27$ $67.49\pm 0.24$ $93.74\pm 0.22$ $64.83\pm 0.46$
LMSPS $\textbf{95.66}\pm\textbf{0.20}$ $\textbf{68.70}\pm\textbf{0.26}$ $\textbf{94.69}\pm\textbf{0.36}$ $\textbf{66.09}\pm\textbf{0.51}$
To answer Q3, we first explore the effectiveness of our search algorithm. In our architecture, our meta-paths are replaced by those meta-paths discovered by other methods. DARTS [30] is the first differentiable search algorithm in neural networks. SPOS [12] is a classic singe-path differentiable algorithm. DARTS and SPOS aim to search operations, like $3\times 3$ convolution, in CNNs. DiffMG [7] and PMMM [25] search for meta-graphs instead of meta-paths. We ignore these differences and focus on the algorithms. The derivation strategies of the four methods are unsuitable for discovering multiple meta-paths. So we changed their derivation strategies to ours to improve their performance. we report the Micro-F1 scores in Table 3. The performance of LMSPS verifies the effectiveness of our search algorithm.
To demonstrate the effectiveness of searched meta-paths, on the one hand, the meta-paths should be effective in the proposed model. On the other hand, the effective meta-paths mainly depend on the dataset instead of the architecture, so the meta-paths should be effective after being generalized to other HGNNs. Because finding meta-paths that work effectively across various HGNNs is a tough task, it has not been achieved by previous works. We verify the generalization of our searched meta-path on the most famous HGNN, HAN [46], and the SOTA metapath-based method SeHGNN [51]. The Micro-F1 scores on three representative datasets are shown in Table 5. After simply replacing the original meta-path set with our searched meta-paths and keeping other settings unchanged, the performance both improve, demonstrating the effectiveness of our searched meta-paths. In addition, we have shown the interpretability of searched meta-paths in Appendix D.
Method DBLP IMDB ACM Freebase
HAN $92.05\pm 0.62$ $64.63\pm 0.58$ $90.79\pm 0.43$ $54.77\pm 1.40$
HAN* $93.54\pm 0.15$ $65.89\pm 0.52$ $92.28\pm 0.47$ $57.13\pm 0.72$
SeHGNN $95.24\pm 0.13$ $68.21\pm 0.32$ $93.87\pm 0.50$ $63.41\pm 0.47$
SeHGNN* $95.57\pm 0.23$ $68.59\pm 0.24$ $94.46\pm 0.18$ $65.37\pm 0.42$
@@ 6.4 Effectiveness of the Search Algorithm and Searched Meta-paths
6.5 Necessity of Long-range Dependency
To answer Q4 and explore the necessity of long-range dependency on heterogeneous graphs, we construct four large-scale datasets with high sparsity based on OGBN-MAG. To avoid inappropriate preference seed settings of randomly removing, we construct fixed heterogeneous graphs by limiting the maximum in-degree related to edge type. Specifically, we gradually reduce the maximum in-degree related to edge type in OGBN-MAG from $50$ to $5$ but leave all nodes unchanged. Details of the four datasets are listed in Appendix A. The test accuracy of LMSPS and SOTA method SeHGNN are shown in Table 5. LMSPS outperforms SeHGNN more significantly with the increasing sparsity. In addition, the leading gap of LMSPS over SeHGNN is more than $4.7\%$ on the highly sparse dataset OGBN-MAG-5. The main difference between SeHGNN and LMSPS is that the former cannot utilize large hops and only use hop $2$ while the latter has a maximum hop of $6$, demonstrating that long-range dependencies are more effective with decreased direct neighbors on heterogeneous graphs.
Table 6: Experiments on small and large datasets to analyze the effects of different blocks in LMSPS. PS means progressive sampling strategy, and SE means sampling evaluation strategy. $\dagger$ means employing all meta-paths and replacing the concatenation operation with the transformer module.
Method DBLP IMDB ACM Freebase OGBN-MAG
Macro-F1 Micro-F1 Macro-F1 Micro-F1 Macro-F1 Micro-F1 Macro-F1 Micro-F1 Test Acc.
LMSPS w/o PS $94.71\pm 0.23$ $95.00\pm 0.19$ $64.85\pm 0.46$ $66.52\pm 0.37$ $93.19\pm 0.34$ $93.14\pm 0.41$ $48.89\pm 0.47$ $61.61\pm 0.51$ $47.66\pm 0.45$
LMSPS w/o SE $95.15\pm 0.28$ $95.48\pm 0.24$ $65.46\pm 0.48$ $67.13\pm 0.47$ $94.20\pm 0.35$ $94.15\pm 0.31$ $52.08\pm 0.33$ $64.84\pm 0.38$ $52.94\pm 0.34$
LMSPS $95.06\pm 0.24$ $95.38\pm 0.21$ $66.85\pm 0.37$ $68.58\pm 0.34$ $94.60\pm 0.42$ $94.57\pm 0.39$ $OOM$ $OOM$ $OOM$
LMSPS $\textbf{95.35}\pm\textbf{0.22}$ $\textbf{95.66}\pm\textbf{0.20}$ $\textbf{66.99}\pm\textbf{0.32}$ $\textbf{68.70}\pm\textbf{0.26}$ $\textbf{94.73}\pm\textbf{0.41}$ $\textbf{94.69}\pm\textbf{0.36}$ $\textbf{53.26}\pm\textbf{0.47}$ $\textbf{66.09}\pm\textbf{0.51}$ $\textbf{54.83}\pm\textbf{0.20}$
6.6 Ablation Study
Two components differentiate our LMSPS from other HGNNs: the search algorithm and semantic fusion without attention. The search algorithm consists of a progressive sampling algorithm and a sampling evaluation strategy. We explore how each of them improves performance through ablation studies under the same settings as the main experiments in Table 1. As shown in Table 6, the performance of LMSPS significantly decreases when removing progressive sampling or sampling evaluation strategy. In addition, employing a transformer block for semantic attention on all meta-paths shows slightly worse performance even if using many more meta-paths and is out-of-memory on larger datasets, indicating that the transformer cannot eliminate the negative effects of noise. It is reasonable because attention values are calculated based on softmax and are positive even for noise.